OpenStudy (anonymous):

When the quadratic equation y = (x − 5)(2x + 3) is written in standard form, what is the value of the coefficient "b"? 7 −7 2 −15

terenzreignz (terenzreignz):

FOIL method is the way to go.

OpenStudy (anonymous):

haven't studied foil is about 5 months and im now doing online lol

terenzreignz (terenzreignz):

\[\huge (px + h)(qx +k)\] time to FOIL F-FIRST TERMS\[\Large (\color{red}{px} + h)(\color{red}{qx} +k)=\color{red}{pqx^2}+...\] O-OUTER TERMS\[\Large (\color{orange}{px} + h)(qx +\color{orange}{k})=pqx^2+\color{orange}{pkx}+...\] I-INNER TERMS \[\Large (px + \color{indigo}h)(\color{indigo}{qx} +k)=pqx^2+pkx+\color{indigo}{qhx}+...\] And finally L-LAST TERMS \[\Large (px + \color{green}h)(qx +\color{green}k)=pqx^2+pkx+qhx+\color{green}{hk}\] And there you have it, the FOIL method \[\LARGE \color{blue}{(px+h)(qx+k)=pqx^2+pkx+qhx+hk}\]

OpenStudy (stamp):

\[(x − 5)(2x + 3)\]\[x(2x+3)-5(2x+3)\]\[2x^2+3x-10x-15\]\[2x^2-7x-15\]Standard form\[ax^2+bx+c\]